# Imaging cross-degree-of-freedom structures in biphoton spatial-polarization hyperentanglement

### v2

@@ 3.1 Certifying spatial-polarization hyperentanglement
Fig. 3 depicts the full spatial maps of the concurrence and biphoton polarization phase in the far field. We calculate an average concurrence of 0.8303 0.0004, which implies strong polarization entanglement with a spatially varying structure across the entire SPDC field. Although biphoton polarization states reside in a two-dimensional Hibert space, the effective dimensionality may be smaller for non-maximally entangled states. Therefore, it is pertinent to estimate a lower bound of the dimensionality in the polarization DoF using the relation [52], where stands for the entanglement of formation [53]. For biphoton polarization states, the entanglement of formation can be derived from concurrence using [51]
where . The average entanglement of formation is then estimated to be . By multiplying the dimensionality lower bounds in spatial and polarization DoFs, we estimate the overall dimensionality of the hyperentangled state to be .
The above result is likely to still be an underestimation of the actual dimensionality since the dimensionality certification protocol has limited resilience against the cross-talk noise [49, 48, 4]. Nevertheless, it is, to the best of our knowledge, the highest dimensionality ever certified for such an entanglement source. Moreover, the varying phase pattern displayed in Fig. 3(b) hints at a sophisticated yet potentially underexplored structure embedded in the high-dimensionality across spatial and polarization DoFs. We note that if one does not resolve the polarization entanglement at each momentum-anticorrelated photon pair, and instead mixes all pairs at a single-pixel detector, the measurements can only reconstruct a lower overall polarization entanglement. Although one can preserve entanglement after mixing by compensating for the phase differences between different pairs, we believe it is more important to understand the cross-DoF structures imposed by the generation process itself, prior to any external engineering, such as spatial compensation.
3.2 Cross-influence between spatial and polarization DoF in SPDC
Mapping the full spatial distribution of polarization entanglement is crucial for exploiting the interplay between spatial and polarization degrees of freedom in SPDC. For instance, having two distinct polarization-entangled states in different biphoton spatial modes allows the creation of photonic cluster states, which is crucial for enabling efficient quantum computation [26, 54].
@@ 3.2 Cross-influence between spatial and polarization DoF in SPDC
The cross-influence between spatial and polarization DoFs in SPDC has two important aspects. First, the exact form of polarization entanglement depends on the transverse momenta of the down-converted photons. As shown in Fig. 3(b), the biphoton polarization phase displays a gradient along the radial direction, indicating polarization entanglement that varies with transverse momentum due to spatial walk-off. This effect occurs because down-converted photons emitted with different transverse momenta experience angle-dependent refractive indices inside the birefringent nonlinear medium and consequently accumulate different phase retardations between orthogonal polarization components. Consequently, a class of highly entangled states of the form of Eq. (1) with spatially varying are generated across the spatial profile of the field. While the spatial profile of polarization entanglement in SPDC has been studied for limited spatial regions or specific sets of spatial modes [1, 40], our results presenting the first full spatial distribution of polarization entanglement in the entire spatial profile of the field, laying the foundation for the controlled generation of desired hyperentangled states.
@@
Phase-matching of the crystal also controls the relation between two-photon transverse momenta and the two-photon polarization phase. As may be noted from Fig. 3(b), the biphoton polarization phase does not vary over the full range of to in the present configuration as a consequence of the near-collinear phase matching condition. In what follows, we shall modify the experimental setup to show that by tuning the crystal orientation, an altered phase matching condition leads to varying over the full parameter space in the spatial profile of the field, which allows access to a much wider class of polarization-entangled states.
@@
We depict our modified setup in Fig. 4(a). In order to capture the enlarged far-field image profile in its entirety, we split the SPDC field into two half-circles using a prism mirror (PM) and again recombine the fields onto the Tpx3Cam sensor. We retain the same far-field imaging scheme, such that the two-photon correlation width remains unchanged from the earlier setup, and we again perform spatially resolved polarization state tomography using the same procedure as discussed previously. In Fig. 4(b),(c), and (d), we depict the far-field intensity Tpx3Cam image, concurrence , and biphoton polarization phase , respectively. We again observe a strong polarization entanglement in the entire spatial extent of the field with an average concurrence of . We notice a marginal increase in the average concurrence compared to the result in Sec. 3.1. This is likely a result of the lower polarization cross-talk since now both signal and idler photons are detected through the transmission port of the PBS. The distribution of again displays a gradient along the radial direction, but this time with having an enlarged parameter space spanning from to . In other words, the setup produces near-maximally entangled states of Eq. (1) with all possible . Thus, our setup can be configured to supply any specific state with a desired value of by post-selecting the corresponding pair of far-field pixels and tuning the phase-matching of the crystal. Such a mechanism can enable encoding spatial mode information in polarization correlations or vice versa. Furthermore, it is also possible to structure to have any desired profile by introducing additional phase differences between H- and V-polarized down-converted photons using spatial light modulators (SLMs) to enable holographic quantum imaging [25] with much higher speed. We note that such phase structures are not limited to only manifesting in momentum space. One can observe similar effects in the near field by exploring different SPDC setup geometries. We may explore near-field phase structuring in future work.
@@
The second important aspect of this cross-influence is between the polarization DoF of the down-converted photons and the spatial DoF of the pump beam. Specifically, the biphoton polarization phase in Eqn. (1) depends on the transverse momentum of the pump beam since [27]. This relation is first theoretically quantified by Li et al. [27]. Here, to experimentally characterize this influence, we introduce a broader angular spectrum in the pump beam by weakly focusing it using a lens with focal length mm placed before the crystal. In Fig. 5, we present the influence of the angular spectrum of the pump beam on the spatial structure of the polarization states. In Fig. 5(a) and (b), we show the far-field intensity profile of the pump and the biphoton momentum correlation profile for the cases of a collimated pump and a focused pump, respectively. In Fig. 5(b), a single superpixel for the signal photons is spatially correlated with multiple superpixels for the idler photons centered around the conjugate superpixel. In other words, after focusing, the increased angular spectrum of the pump beam substantially widens the biphoton correlation width. In Fig. 5(c) and (e), we depict the results for polarization state tomography for the case of the collimated pump. To illustrate the influence of the transverse momentum of the pump beam on the polarization entanglement of the down-converted photons, we map the concurrence and phase onto pixel coordinates of the joint momentum of signal and idler, which is essentially the transverse momentum of the pump beam . Here, the polarization entanglement is found in a few superpixels within the narrower correlation profile, and displays little variation within the narrow angular spectrum of the collimated pump. In Fig. 5(d) and (f), we depict the results for polarization state tomography for the case of the focused pump. The concurrence map indicates the presence of polarization entanglement in all superpixels within the widened correlation profile, with an average concurrence of . The decrease in concurrence upon focusing the pump likely results from lower count rates and higher statistical fluctuations in individual pixels. In contrast to Fig. 5(e), the phase map illustrates a gradient of dependent on the transverse momentum of the pump beam, which is in good agreement with the theoretical predictions in Fig. 2(a) of [27]. To the best of our knowledge, this is the first direct measurement of the cross-influence between the spatial DoF of the pump beam and the polarization DoF of the down-converted photons. Our work opens up the opportunity to control the spatial structure of polarization states by jointly manipulating the polarization and spatial modes of the pump, which has implications in the study of topological structures of quantum light [55, 56, 57]. For instance, one can pump a paired crystal with a beam with spatially structured polarization and directly shape the correlation between spatial mode and polarization in the nonlocal optical skyrmions[24].
4 Conclusion and Outlook
@@
In this work, we image the cross-DoF structure in spatial-polarization hyperentangled biphotons produced from SPDC using a data-driven camera capable of time-stamping single photons. We first demonstrate the rapid characterization of hyperentanglement with a cumulative measurement time of only 17 minutes. In particular, we certify the entanglement dimensionality to be about 148 in the spatial DoF using the method discussed in Refs. [48, 49]. By performing spatially-resolved polarization state tomography, we confirm the generation of an entire class of near-maximally polarization-entangled states with an average concurrence of 0.83030.0004 in the entire spatial profile of the SPDC field, thereby certifying a spatial-polarization hyperentanglement with an estimated total dimensionality of 251. We then present the cross-influence between spatial and polarization DoF. By producing the first complete spatial map of the polarization state generated from SPDC, we illustrate the dependence of polarization entanglement on the transverse momenta of the down-converted photons in the entire spatial profile of the SPDC field. Using a weakly focused pump beam, we, for the first time, experimentally characterize the cross-influence between the polarization DoF of the down-converted photons and the spatial DoF of the pump beam.
@@
Our results could be important for future work aimed at harnessing hyperentangled states for high-dimensional quantum information processing applications. For instance, time-stamping single photons with a data-driven camera could significantly reduce the acquisition time requirements in polarization entanglement-enabled holography [25], potentially enabling fast super-resolution imaging and microscopy [58]. Our results on cross-influence between different DoFs could extend the recent demonstration of hiding images in quantum correlations [29] to hiding and rapidly reconstructing complex phase images in spatial and polarization correlations. On the fundamental side, our work can be extended to explore the spatial-polarization cross-influence in different hyperentanglement generation schemes. For instance, it is possible to spatially resolve the polarization entanglement produced from post-selecting down-converted photons from a single Type-II nonlinear crystal [59, 28], thereby further deepening our understanding of the cross-DoF influence in SPDC. It may also be possible to build on existing work on exchange phases in Hong-Ou-Mandel interference involving high-dimensional hyperentangled photons [60] and spatially resolve such effects for a hyperentangled state. Although the current Tpx3Cam system has a low overall photon detection efficiency of 8 % [41] and is thus not yet suited for studies involving Bell-type tests of nonlocality [61], its applicability could see further expansion by implementing an image intensifier with higher quantum efficiency and higher gain. For instance, a minimum detection efficiency of is required to close the fair-sampling loophole in Bell tests [62].
Funding National Science Foundation (2244031); U.S. Department of Energy (76295); Natural Sciences and Engineering Research Council of Canada (RGPIN/2017-06880); Canada First Research Excellence Fund (072623).
Acknowledgement The authors thank X. Gao, A. D’Errico, B. Braverman, S. Karan, and M. Krenn for fruitful discussions. C. Li acknowledges P. Svihra for helpful advice on data processing.
Data availability All relevant data are available from the corresponding author upon reasonable request.
Disclosures The authors declare no conflict of interest.

### v3

@@ III.2 Certifying total attainable dimensionality of in the joint spatial-polarization bases
Although two-photon polarization states reside in a two-dimensional Hilbert space, the effective dimensionality may be smaller for non-maximally entangled states. Therefore, it is pertinent to estimate a lower bound of the dimensionality in the polarization DoF using the relation [37], where stands for the entanglement of formation [50]. For two-photon polarization states, the entanglement of formation can be derived from concurrence using [49]
where . The average entanglement of formation is then estimated to be . By multiplying the dimensionality lower bounds in spatial and polarization DoFs, we estimate the total attainable dimensionality of the entangled state to be . We note that this result is likely to still be an underestimation of the actual dimensionality since the dimensionality certification protocol has limited resilience against the cross-talk noise [14, 3, 16], a more accurate estimation may be explored in a future work.
III.3 Native spatial-polarization coupling in SPDC processes
@@
The spatial-polarization coupling in SPDC has two important aspects. First, the exact form of polarization entanglement depends on the transverse momenta of the down-converted photons. This effect, which is displayed in Fig. 2(b), occurs because down-converted photons emitted with different transverse momenta experience angle-dependent refractive indices inside the birefringent nonlinear medium and consequently accumulate different phase retardations between orthogonal polarization components. Consequently, a class of highly entangled states of the form of Eq. (1) with spatially varying is generated across the spatial profile of the field. While the spatial profile of polarization entanglement in SPDC has been studied for limited spatial regions or specific sets of spatial modes [2, 1], our results present the first full spatial distribution of polarization entanglement in the entire spatial profile of the field, laying the foundation for the controlled generation of desired photonic entangled states.
@@
Phase-matching of the crystal also controls the relation between two-photon transverse momenta and the two-photon polarization phase. As may be noted from Fig. 2(b), the two-photon polarization phase does not vary over the full range of to in the present configuration as a consequence of the near-collinear phase matching condition. In what follows, we modify the experimental setup to show that by tuning the crystal orientation, an altered phase matching condition leads to varying over the full parameter space in the spatial profile of the field, which allows access to a much wider class of polarization-entangled states.
@@
We depict our modified setup in Fig. 4(a). In order to capture the enlarged far-field image profile in its entirety, we deterministically split the SPDC field into two half-circles using a prism mirror (PM) and recombine the fields onto the Tpx3Cam sensor. We retain the same far-field imaging scheme, such that the two-photon correlation width remains unchanged from the earlier setup, and we again perform spatially resolved polarization state tomography using the same procedure as discussed previously. In Fig. 4(b),(c), and (d), we depict the far-field intensity Tpx3Cam image, concurrence , and two-photon polarization phase , respectively. We again observe a strong polarization entanglement in the entire spatial extent of the field with an average concurrence of . We notice a marginal increase in the average concurrence compared to the result in Sec. III.2. This is likely a result of the lower polarization cross-talk since now both signal and idler photons are detected through the transmission port of the PBS. The distribution of again displays a gradient along the radial direction, but this time with having an enlarged parameter space spanning from to . In other words, the setup produces near-maximally entangled states of Eq. (1) with all possible . Thus, our setup can be configured to supply any specific state with a desired value of by post-selecting the corresponding pair of far-field pixels and tuning the phase-matching of the crystal. Such a mechanism can enable encoding spatial mode information in polarization correlations or vice versa, which could have important implications for quantum key distribution protocols using qudit-like states [43]. Furthermore, it is also possible to structure to have any desired profile by introducing additional phase differences between H- and V-polarized down-converted photons using SLMs to enable holographic quantum imaging [9] with much higher speed. We note that such phase structures are not limited to only manifesting in momentum space. One can observe similar effects in the near field by exploring different SPDC setup geometries. We may explore near-field phase structuring in future work.
@@
The second important aspect of spatial-polarization structure is between the polarization DoF of the down-converted photons and the spatial DoF of the pump beam. Specifically, the two-photon polarization phase in Eqn. (1) depends on the transverse momentum of the pump beam since [30]. This relation is first theoretically quantified by Li et al. [30]. Here, to experimentally characterize this influence, we introduce a broader angular spectrum in the pump beam by weakly focusing it using a lens with focal length mm placed before the crystal. In Fig. 5, we present the influence of the angular spectrum of the pump beam on the spatial structure of the polarization states. In Fig. 5(a) and (b), we show the far-field intensity profile of the pump and the two-photon momentum correlation profile for the cases of a collimated pump and a focused pump, respectively. In Fig. 5(b), a single superpixel for the signal photons is spatially correlated with multiple superpixels for the idler photons centered around the conjugate superpixel. In other words, after focusing, the increased angular spectrum of the pump beam substantially widens the two-photon correlation width. In Fig. 5(c) and (e), we depict the results for polarization state tomography for the case of the collimated pump. To illustrate the influence of the transverse momentum of the pump beam on the polarization entanglement of the down-converted photons, we map the concurrence and phase onto pixel coordinates of the joint momentum of signal and idler, which is essentially the transverse momentum of the pump beam . Here, the polarization entanglement is found in a few superpixels within the narrower correlation profile, and displays little variation within the narrow angular spectrum of the collimated pump. In Fig. 5(d) and (f), we depict the results for polarization state tomography for the case of the focused pump. The concurrence map indicates the presence of polarization entanglement in all superpixels within the widened correlation profile, with an average concurrence of . The decrease in concurrence upon focusing the pump likely results from lower count rates and higher statistical fluctuations in individual pixels. In contrast to Fig. 5(e), the phase map illustrates a gradient of dependent on the transverse momentum of the pump beam, which is in good agreement with the theoretical predictions in Fig. 2(a) of [30]. To the best of our knowledge, this is the first direct measurement of the cross-influence between the spatial DoF of the pump beam and the polarization DoF of the down-converted photons. Our work opens up the opportunity to control the spatial structure of polarization states by jointly manipulating the polarization and spatial modes of the pump, which has implications in the study of topological structures of quantum light [21, 44, 34]. For instance, one can pump a paired crystal with a beam with spatially structured polarization and directly shape the correlation between spatial mode and polarization in the nonlocal optical skyrmions[40].
IV Conclusion and Outlook
@@
In this work, we image the spatially varying polarization entanglement in a two-photon state produced from SPDC using a data-driven camera capable of time-stamping single photons. In contrast to earlier studies that have imposed spatial-polarization structures by modulating a postselected subset of the full SPDC field, our work reveals those that are inherent to the SPDC process. By performing spatially-resolved polarization state tomography, we confirm the generation of an entire class of near-maximally polarization-entangled states with an average concurrence of 0.83030.0004 in the entire spatial profile of the SPDC field. Additionally, we certify an attainable entanglement dimensionality of 148 in the spatial DoF using the method discussed in Refs. [3, 14], which indicates a total attainable dimensionality of 251 in the joint spatial-polarization basis. We then present the experimentally characterized spatial-polarization structure. By producing the first complete spatial map of the polarization state generated from SPDC, we illustrate the dependence of polarization entanglement on the transverse momenta of the down-converted photons in the entire spatial profile of the SPDC field. Using a weakly focused pump beam, we, for the first time, experimentally characterize the cross-influence between the polarization DoF of the down-converted photons and the spatial DoF of the pump beam. Although the current Tpx3Cam system has a low overall photon detection efficiency of 8 % [48] and is thus not yet suited for studies involving Bell-type tests of nonlocality [52], its applicability could see further expansion by implementing an image intensifier with higher quantum efficiency and higher gain. For instance, a minimum detection efficiency of is required to close the fair-sampling loophole in Bell tests [19].
@@
Our results could be important for future work aimed at harnessing SPDC-based entanglement generation for high-dimensional quantum information processing applications [13, 16]. For instance, time-stamping single photons with a data-driven camera could significantly reduce the acquisition time requirements in polarization entanglement-enabled holography [9], potentially enabling fast super-resolution imaging and microscopy [8]. Our results on cross-influence between different DoFs could extend the recent demonstration of hiding images in quantum correlations [47] to hiding and rapidly reconstructing complex phase images in spatial and polarization correlations. On the fundamental side, our work can be extended to explore the spatial-polarization coupling in different entanglement generation schemes. For instance, it is possible to spatially resolve the polarization entanglement produced from post-selecting down-converted photons from a single Type-II nonlinear crystal [26, 53], thereby further deepening our understanding of the cross-DoF coupling in SPDC. Furthermore, our work could have important implications for studying and engineering hyperentangled photon states for various applications, such as noise-resilient quantum illumination [41], quantum key distribution with high secure key rates [51, 25, 36]. It may also be possible to build on existing work on exchange phases in Hong-Ou-Mandel interference involving high-dimensional hyperentangled photons [31] and spatially resolve such effects for a hyperentangled state.
Acknowledgements.
The authors thank X. Gao, A. D’Errico, B. Braverman, S. Karan, and M. Krenn for fruitful discussions. C. L. acknowledges P. Svihra for helpful advice on data processing. The portion of the work performed at the University of Ottawa was supported by the Canada Research Chairs program under Award 950-231657, the Natural Sciences and Engineering Research Council of Canada under Alliance Consortia Quantum Grant ALLRP 578468 - 22, Discovery Grant RGPIN/2017-06880, and the Canada First Research Excellence Fund Award 072623. In addition, R.W.B. acknowledges support through U.S. National Science Foundation Award No. 2138174 and U.S. Department of Energy Award No. FWP 76295.Appendix A Observation of spatial entanglement through the violation of Einstein-Podolsky-Rosen criteria
In Fig. 6, we depict the experimentally measured - and -components of the spatial correlation profile in both the position and momentum basis. We note the strong position correlations and momentum anti-correlations between the photons, which is a characteristic feature of spatial entanglement. Specifically, Fig. 6(c) and (f) depict the biphoton correlation profiles in the joint momentum and position coordinates of the down-converted photons. From these results, we conclude that in the Tpx3Cam sensor plane, individual discrete momentum and position modes are best approximated by -pixel large superpixels.
The results also allow us to quantitatively verify spatial entanglement by demonstrating violations of the Einstein-Podolsky-Rosen (EPR) criteria [42, 4]. We first obtain near-field (NF) and far-field (FF) correlation widths and , respectively, at the Tpx3Cam sensor plane by fitting Gaussians to the spatial correlation profiles of Fig. 6 in the - and -directions [32, 29, 4]. We then calculate the position and momentum uncertainties using the relations
where and stand for the momentum and position of the signal (idler) photons, respectively, represent the x- and y-components of the quantities, nm-1 is the wavevector of the signal and idler photons, mm is the effective focal length of our far-field imaging system and is the magnification of our near-field imaging system. We summarize our results in Table 1 and compute the conditional Heisenberg uncertainty products as
which clearly violates the EPR criteria in both the x- and y-directions. We note that the asymmetry observed in the x-components of the position correlation, which is depicted in Fig. 6(d), is likely a consequence of a distortion in the pump’s intensity profile. As depicted in Fig. 1(c) in the main text, the near-field singles rates display a non-Gaussian spatial profile as a result of the distorted pump profile. Although this imperfection may affect the verification of spatial entanglement using the EPR criterion, which requires the pump beam to have a Gaussian profile [29], it does not refute our conclusions regarding the certification of hyperentanglement, as our entanglement certification protocol does not rely on Gaussian approximations of the SPDC field profiles [14].
| Quantity | Values | Units |
| m | ||
| m | ||
| m | ||
| m |
Appendix B Estimation of fidelity lower bound
Following the approaches described in [35], we certify the dimensionality of entanglement in the spatial degree of freedom using correlations in two mutually unbiased bases (MUBs) [14, 3]. In the main text, we have defined the two MUBs to be the discrete momentum basis and the discrete position basis .
The fidelity of the experimentally measured state to a maximally entangled state is defined as:
where
Since for an arbitrary state with a Schmidt number of , the fidelity satisfies
the entanglement dimensionality of a state with must be at least . As we will show in the following, correlations in the two MUBs will allow us to obtain a lower bound for the fidelity via
The coincidence counts measured in the discrete momentum basis allow us to calculate using
Supplementing the coincidence counts measured in the discrete position basis allows us to calculate , the lower bound of , via
where